Big-Pear Corp. is considering replacing its existing equipment that is used to produce smart cell phones. This existing equipment was purchase 3 years ago at a base price of $40,000. Installation costs at the time for the machine were $8,000. The existing equipment is considered a 5-year class for MACRS. The existing equipment can be sold today for $40,000 and for $30,000 in 3 years. The new equipment has a purchase price of $140,000 and is also considered a 5-year class for MACRS. Installation costs for the new equipment are $7,000. The estimated salvage value of the new equipment is $80,000. This new equipment is more efficient than the existing one and thus savings before taxes using the new equipment are $18,000 a year. Due to these savings, inventories will see a one time reduction of $2,000 at the time of replacement. The company's marginal tax rate is 20% and the cost of capital is 12%. For this project, what is the incremental cash flow in year 2?

MACRS Fixed Annual Expense Percentages by Recovery Class
Year 3-Year 5-Year 7-Year 10-Year 15-Year
1 33.33% 20.00% 14.29% 10.00% 5.00%
2 44.45% 32.00% 24.49% 18.00% 9.50%
3 14.81% 19.20% 17.49% 14.40% 8.55%
4 7.41% 11.52% 12.49% 11.52% 7.70%
5 11.52% 8.93% 9.22% 6.93%
6 5.76% 8.93% 7.37% 6.23%
7 8.93% 6.55% 5.90%
8 4.45% 6.55% 5.90%
9 6.56% 5.91%
10 6.55% 5.90%
11 3.28% 5.91%
12 5.90%
13 5.91%
14 5.90%
15 5.91%
16 2.95%

For your answer, round to the nearest dollar, do not enter the $ sign, use commas to separate thousands, use a negative sign in front of first number is the cash flow is negative (do not use parenthesis to indicate negative cash flows). For example, if your answer is $3,005.87 then enter 3,006; if your answer is -$1,200.25 then enter-1,200

For this project, the incremental cash flow in year 2 is:
Answer #1

Answer :

Calculation of Incremental Cash Flows for year 2:

Incremental Cash Flow = Before tax Savings * (1 - tax rate) + Tax shield on Incremental Depreciation

Tax shield on Incremental Depreciation = (Depreciation on new machine - Depreciation on Old machine) * Tax rate

Depreciation on New Machine = (Base Price + Installation) * Depreciation rate for 2nd year

= (140,000 + 7000) * 32%

= 47040

Depreciation on Existing Machine = (Base Price + Installation) * Depreciation rate for 5th year

= (40000 + 8000) * 11.52%

= 5529.6

Note : For existing machine rate of year 5 will be taken as it was purchased three years ago and we are calculating incremental cash flows for year 2

Tax shield on Incremental Depreciation = (47040 - 5529.6) * 20%

= 8302.08

Incremental Cash Flow = 18000 * (1 - 0.20) + 8302.08

= 22702.08 or 22,702
Answer #2

Answer :

Calculation of Incremental Cash Flows for year 2:

Incremental Cash Flow = Before tax Savings * (1 - tax rate) + Tax shield on Incremental Depreciation

Tax shield on Incremental Depreciation = (Depreciation on new machine - Depreciation on Old machine) * Tax rate

Depreciation on New Machine = (Base Price + Installation) * Depreciation rate for 2nd year

= (140,000 + 7000) * 32%

= 47040

Depreciation on Existing Machine = (Base Price + Installation) * Depreciation rate for 5th year

= (40000 + 8000) * 11.52%

= 5529.6

Note : For existing machine rate of year 5 will be taken as it was purchased three years ago and we are calculating incremental cash flows for year 2

Tax shield on Incremental Depreciation = (47040 - 5529.6) * 20%

= 8302.08

Incremental Cash Flow = 18000 * (1 - 0.20) + 8302.08

= 22702.08 or 22,702


answered by: ANURANJAN SARSAM